# Exact Diagonalization, Matrix Product States and Conformal Perturbation Theory Study of a 3D Ising Fuzzy Sphere Model

### v1

@@ B.4 Matrix elements for shifts of $T$ and $C$ from ${\cal V}=\varepsilon$
The conformal bootstrap study Chang:2024whx , which included the stress tensor among external states, determined:
@@
We will use this extremely accurate value. A less accurate prior conformal bootstrap determination was $0.958(7)$ Poland:2023bny , from the five-point bootstrap. To extract the energy level shift one has to take the limit of the three-point function (Chang:2024whx, , Eq. (2.2)) putting the stress tensors at 0 and $\infty$, and integrate the scalar over the sphere. This gives
@@
We note that the quantity called the “$f_{TT\varepsilon}$ OPE coefficient” in the fuzzy bootstrap study Hu:2023xak has exactly the same normalization as $f^{\rm shift}_{TT\varepsilon}$ and should agree with it in the $N\to\infty$ limit. Yet they reported the value 0.8658(69), in discrepancy with CB. This is likely due to the difficulties of extrapolating to infinite volume in their setup. In Sec. 6 we present a different way of extracting $f^{\rm shift}_{TT\varepsilon}$ from the fuzzy sphere data, which is less affected by finite size corrections and gives a value in agreement with CB.
@@
The coefficients $\lambda_{C\varepsilon C}^{i}$ have been determined in a five-point bootstrap study Poland:2023bny (improving on Poland:2023vpn ). Using the last column in (Poland:2023bny, , Table 3):
@@
one obtains $f_{CC\varepsilon}^{\rm shift}=0.40(23)$. [^25] Since the correlation matrix of the $1\sigma$ errors on $\lambda_{C\varepsilon C}^{i}$ is not reported in Poland:2023bny , here we just combined all the errors with appropriate coefficients and $\pm$ signs chosen to maximize the overall error on $f_{CC\varepsilon}^{\rm shift}$. Recently, the authors of Poland:2023bny repeated their analysis including the high precision value of $\lambda^{0}_{TT\varepsilon}$ from Chang:2024whx and extracting directly the linear combination (91) of interest to us. This gives a more accurate determination petar :
@@
Appendix C CPT results of Reinicke Reinicke1987a ; Reinicke1987b
Reinicke Reinicke1987a ; Reinicke1987b developed CPT for the 2D Ising CFT. He considered Ising CFT states on $S^{1}$ of radius $R=1$:
$|h,\bar{h}\rangle=|X+r,\bar{X}+\bar{r}\rangle,\quad r,\bar{r}\in\mathbb{Z}_{\geqslant 0},$ (94)
corresponding to local operators $\Phi_{h,\bar{h}}$ of conformal weights $h=X+r,\bar{h}=\bar{X}+\bar{r}$. The Virasoro descendants of $\mathds{1},\varepsilon,\sigma$ at level $r,\bar{r}$ where ($X,\bar{X}$)$\in\{(0,0),(1/2,1/2),(1/16,1/16)\}$ are the weights of the Virasoro primary. [^26] In Reinicke’s papers $X,\bar{X}$ are denoted $\Delta,\bar{\Delta}$. (Some of these descendant levels do not exist since they are null, i.e., have zero norm.) Their unperturbed energies are $E_{h,\bar{h}}=h+\bar{h}-c/12$, and the energy gaps are
@@ Appendix C CPT results of Reinicke Reinicke1987a ; Reinicke1987b
[…]
where $\mathcal{V}$ is one of the three operators $\varepsilon$, $T\bar{T}$ or $T^{2}+\bar{T}^{2}$ (see footnote 3). He computed the energy gaps $\mathcal{F}_{h,\bar{h}}$ in perturbation theory to fourth order in $G_{\varepsilon}$ Reinicke1987a , and to first order in the other two couplings Reinicke1987b , for the cases when the unperturbed energy level $|h,\bar{h}\rangle$ has degeneracy exactly one, which happens for the 2D Ising CFT for
@@ Appendix D Ising Hamiltonian on the fuzzy sphere
The fuzzy sphere formulation of the 2D Ising model is based on the study of quantum Hall models on the sphere (Haldane1983, ; Haldane1985, ; Fano1986, ). We consider electrons living on a sphere with a $4\pi s$ monopole at the center. Each electron has spin-$1/2$ ($\sigma=\mathbin{\uparrow}\downarrow$) and orbital momentum $m=-s,-s+1,...,s$ degrees of freedom, where the latter stems from the $12s+$ 1-fold degeneracy of each lowest landau level (LLL) orbital in this model. Practically, $s$ determines the resolution of the simulation since $N=2s+1$ is the total number of orbitals per spin orientation. We consider a half-filled configuration where the number of electrons $N_{e}$ is equal to $N$.
Following the original construction presented in Ref. Zhu2023 , the Hamiltonian of the (2+1)D transverse field Ising model on the fuzzy sphere is given by
@@
$H_{00}$, up to normal ordering, is the well-studied quantum Hall Hamiltonian. The ground state for the chosen $V_{0}\gg V_{1}$ at $N_{e}=N$ electrons is a Mott insulator whose spin degrees of freedom form a ferromagnet (Sondhi1993, ; Girvin2000, ). $H_{xx}$ breaks the SU($2$) invariance down to $U(1)$ and the ferromagnet polarizes in the $x$-direction. Finally, $H_{t,z}$ is a polarizing field in the transverse direction and competes with $H_{xx}$. In the limit $h\rightarrow\infty$, the spin degrees of freedom form a paramagnet in the $z$-direction. An Ising transition is therefore expected at intermediate $h$ as was shown in Ref. Zhu2023 . Note that we performed a rotation of the spin-$1/2$ degree of freedom around $\sigma^{y}$ compared to the original Hamiltonian.
@@ E.1 Exact diagonalization
$\mathcal{H}_{N_{e},\bar{m},p}=\mathrm{Span}\Bigg{\{}{\prod_{i=1}^{N_{e}}c^{\dagger}_{m_{i},\sigma_{i}}\ket{0}}\Bigg{\}},$ (114)
@@
In order to find the low-energy spectrum of $H$ from Eq. (112) in the Hilbert space of a given symmetry sector, including low-lying excitations, we make use of Arnoldi iteration implemented in the ARPACK Lehoucq1998arpack package. Similar to Lanczos iteration, the algorithm requires nothing but the implementation of the matrix-vector product |$\ket{\phi}=H\ket{\psi}$ to converge a given number of extremal (in this case minimal) eigenvalues. Unfortunately, one of the limiting aspects in terms of reachable Hilbert space dimensions is the integer data type used in the available ARPACK implementation, where the 32-bit nature limited the available system sizes in our simulations. So while $N=20$ ED simulations are technically feasible with a dimension of about 1.5 billion, because of ARPACK issues the diagonalizations do not seem to converge. Nevertheless, systems sizes of up to $N=18$ with corresponding maximal (symmetry resolved) Hilbert space dimensions of 113 million could be successfully studied.
@@ E.2 ITensor MPS Implementation
We use ITensors.jl (ITensor, ; ITensor-r0.3, ) to perform our matrix product state computations. Matrix product state computations for fractional Quantum Hall systems have been thoroughly explored (Shibata2001, ; Feiguin2008, ; Kovrizhin2010, ; ZaleteliDMRG1, ; ZaleteliDMRG2, ) . The fuzzy sphere is represented by a chain of $2\times N$ spinless electrons, in ascending order of orbital momentum. The $2$ flavor of spins (spin up and down) at a given momentum correspond to consecutive sites. We implemented the U(1) conservation of the electronic charge and the orbital momentum, as well as conservation of the spin parity along the $z$ direction.
@@
The main algorithm we used to obtain the ground states is two-site DMRG (White1992, ). Due to the large number of symmetries, and in particular the interplay between charge and momentum conservation, two-site updates are not sufficient to ensure the ergodicity of the variational optimization. In fact, if the starting state is a simple local product state in the symmetric basis, it will remain a product state. To remedy to this issue, we initialize our state in an inversion-symmetric product state in the $L^{z}=0$ sector. We then apply the Hamiltonian to the product state, and truncate the resulting MPS to a low bond-dimension, before using ladder operators to build an initial state in the targeted symmetry sector. Finally, at each bond dimension $\chi$, we first perform noisy DMRG (White2005, ) sweeps to explore phase space followed by conventional noiseless DMRG to converge to the true ground state. The amplitude of the noise was generally kept at $10^{-5}$. Our convergence criterion at a given $\chi$ are variations of the energy and the mid-chain entropy below $10^{-7}$.

### v2

@@ B.4 Matrix elements for shifts of $T$ and $C$ from ${\cal V}=\varepsilon$
The conformal bootstrap study [45], which included the stress tensor among external states, determined:
@@
We will use this extremely accurate value. A less accurate prior conformal bootstrap determination was $0.958(7)$ [70], from the five-point bootstrap. To extract the energy level shift one has to take the limit of the three-point function [45, Eq. (2.2)] putting the stress tensors at 0 and $\infty$, and integrate the scalar over the sphere. This gives
@@
We note that the quantity called the $f_{TT\varepsilon}$ OPE coefficient” in the fuzzy bootstrap study [14] has exactly the same normalization as $f^{\rm shift}_{TT\varepsilon}$ and should agree with it in the $N\to\infty$ limit. Yet they reported the value 0.8658(69), in discrepancy with CB. This is likely due to the difficulties of extrapolating to infinite volume in their setup. In Sec. 6 we present a different way of extracting $f^{\rm shift}_{TT\varepsilon}$ from the fuzzy sphere data, which is less affected by finite size corrections and gives a value in agreement with CB.
@@
The coefficients $\lambda_{C\varepsilon C}^{i}$ have been determined in a five-point bootstrap study [70] (improving on [69]). Using the last column in [70, Table 3]:
@@
one obtains $f_{CC\varepsilon}^{\rm shift}=0.40(23)$. [^29] Since the correlation matrix of the $1\sigma$ errors on $\lambda_{C\varepsilon C}^{i}$ is not reported in [70], here we just combined all the errors with appropriate coefficients and $\pm$ signs chosen to maximize the overall error on $f_{CC\varepsilon}^{\rm shift}$. Recently, the authors of [70] repeated their analysis including the high precision value of $\lambda^{0}_{TT\varepsilon}$ from [45] and extracting directly the linear combination (91) of interest to us. This gives a more accurate determination [67]:
@@
Appendix C CPT results of Reinicke [32, 33]
Reinicke [32, 33] developed CPT for the 2D Ising CFT. He considered Ising CFT states on $S^{1}$ of radius $R=1$:
[…]
@@ Appendix C CPT results of Reinicke [32, 33]
$V=\sum_{\mathcal{V}}G_{\mathcal{V}}\int_{S^{1}}dx\,\mathcal{V}(0,x),$ (96)
where $\mathcal{V}$ is one of the three operators $\varepsilon$, $T\bar{T}$ or $T^{2}+\bar{T}^{2}$ (see footnote 3). He computed the energy gaps $\mathcal{F}_{h,\bar{h}}$ in perturbation theory to fourth order in $G_{\varepsilon}$ [32], and to first order in the other two couplings [33], for the cases when the unperturbed energy level $|h,\bar{h}\rangle$ has degeneracy exactly one, which happens for the 2D Ising CFT for
@@ Appendix D Ising Hamiltonian on the fuzzy sphere
The fuzzy sphere formulation of the 2D Ising model is based on the study of quantum Hall models on the sphere [75, 76, 77]. We consider electrons living on a sphere with a $4\pi s$ monopole at the center. Each electron has spin-$1/2$ ($\sigma=\mathbin{\uparrow}\hskip-5.0pt\downarrow$) and orbital momentum m=s,s+1,,sm=-s,-s+1,...,s degrees of freedom, where the latter stems from the $2s+1$-fold degeneracy of each lowest landau level (LLL) orbital in this model. Practically, $s$ determines the resolution of the simulation since $N=2s+1$ is the total number of orbitals per spin orientation. We consider a half-filled configuration where the number of electrons $N_{e}$ is equal to $N$.
Following the original construction presented in Ref. [1], the Hamiltonian of the (2+1)D transverse field Ising model on the fuzzy sphere is given by
@@
$H_{00}$, up to normal ordering, is the well-studied quantum Hall Hamiltonian. The ground state for the chosen $V_{0}\gg V_{1}$ at $N_{e}=N$ electrons is a Mott insulator whose spin degrees of freedom form a ferromagnet [79, 80]. $H_{xx}$ breaks the SU($2$) invariance down to $U(1)$ and the ferromagnet polarizes in the $x$-direction. Finally, $H_{t,z}$ is a polarizing field in the transverse direction and competes with $H_{xx}$. In the limit $h\rightarrow\infty$, the spin degrees of freedom form a paramagnet in the $z$-direction. An Ising transition is therefore expected at intermediate $h$ as was shown in Ref. [1]. Note that we performed a rotation of the spin-$1/2$ degree of freedom around $\sigma^{y}$ compared to the original Hamiltonian.
@@ E.1 Exact diagonalization
[…]
@@
[…]
@@ E.2 ITensor MPS Implementation
We use ITensors.jl [82, 83] to perform our matrix product state computations. Matrix product state computations for fractional Quantum Hall systems have been thoroughly explored [84, 85, 86, 87, 88] . The fuzzy sphere is represented by a chain of $2\times N$ spinless electrons, in ascending order of orbital momentum. The $2$ flavor of spins (spin up and down) at a given momentum correspond to consecutive sites. We implemented the U(1) conservation of the electronic charge and the orbital momentum, as well as conservation of the spin parity along the $z$ direction.
@@
The main algorithm we used to obtain the ground states is two-site DMRG [89]. Due to the large number of symmetries, and in particular the interplay between charge and momentum conservation, two-site updates are not sufficient to ensure the ergodicity of the variational optimization. In fact, if the starting state is a simple local product state in the symmetric basis, it will remain a product state. To remedy to this issue, we initialize our state in an inversion-symmetric product state in the $L^{z}=0$ sector. We then apply the Hamiltonian to the product state, and truncate the resulting MPS to a low bond-dimension, before using ladder operators to build an initial state in the targeted symmetry sector. Finally, at each bond dimension $\chi$, we first perform noisy DMRG [90] sweeps to explore phase space followed by conventional noiseless DMRG to converge to the true ground state. The amplitude of the noise was generally kept at $10^{-5}$. Our convergence criterion at a given $\chi$ are variations of the energy and the mid-chain entropy below $10^{-7}$.